# Patrick Cheridito and Moritz Weiss

### v1

@@ 3.4 Benchmark algorithms
We compare our trained reinforcement learning algorithm with two heuristic benchmarks and an alternative reinforcement learning algorithm using Dirichlet distributions. The Dirichlet-RL algorithm is explained in Appendix C. The heuristic strategies trade according to the following rules:
@@
  • Time-weighed average price [^4] This is abbreviated with TWAP and represents a standard execution algorithm banks and hedge funds use. algorithm (TWAP algorithm): Divides the position into blocks of size $M/N.$ For $n\in\{0,1,\dots,N-1\},$ it posts a limit sell order of size $M/N$ at the best ask price every time step $t_{n}.$
@@ 4.1 Actor-critic policy gradient
The trade execution problem described in Section 3 is a stochastic control problem. The transition probabilities from state $s_{n}$ to $s_{n+1}$ and the reward function $r(\cdot)$ are not available in closed form, but we can simulate them. Therefore, the problem can be solved with reinforcement learning. Actor-critic policy gradient algorithms are a class of reinforcement learning algorithms that aim to find an optimal stochastic policy $\pi_{\theta}(\cdot\mid s_{n}),$ referred to as actor, by estimating the gradient of the objective function $J(\pi_{\theta}).$ A separate parameterized value or advantage function, referred to as critic, is used to reduce the variance of the gradient estimate. We use a neural network with weights $\theta$ to map states $s_{n}$ to the parameters of the policy’s density. The gradient of the objective function can be computed as
$\nabla_{\theta}J(\pi_{\theta})=\underset{a\sim\pi_{\theta}}{\operatorname{\mathbb{E}}}\left[\sum_{n=0}^{N-1}\nabla_{\theta}\log\pi_{\theta}(a_{n}\mid s_{n})A^{\pi_{\theta}}(s_{n},a_{n})\right],$ (4)
by the policy gradient theorem; see e.g. Schulman et al. (2015). We note that the actions $a_{0},\dots a_{N-1}$ are sampled from a conditional distribution with density $\pi_{\theta}(\cdot\mid s_{n})$ and the states $(s_{0},\dots,s_{N-1})$ are sampled from the market environment. Here, $A^{\pi_{\theta}}$ is the advantage function, which is defined as
$A^{\pi_{\theta}}(s,a)=\underset{a\sim\pi_{\theta}}{\operatorname{\mathbb{E}}}\left[\sum_{l=n}^{N-1}r(s_{l},a_{l})\mid s_{n}=s,a_{n}=a\right]-\underset{a\sim\pi_{\theta}}{\operatorname{\mathbb{E}}}\left[\sum_{l=n}^{N-1}r(s_{l},a_{l})\mid s_{n}=s\right],$ (5)
@@ 4.2 Policy gradient with logistic-normal distribution
In most applications of policy gradients with continuous action spaces, the multivariate normal distribution is used as an action distribution. However, in our application, the actions must be contained in the simplex. Therefore, we require a distribution with support on the simplex. The works Tian et al. (2022) and Winkel et al. (2023) have used the Dirichlet distribution to achieve this. But we found that the Dirichlet distribution performs sub-optimally in our experiments, even with extensive hyperparameter tuning. An alternative to the Dirichlet distribution is the logistic-normal distribution, which is also defined on the simplex.
@@
$\pi(a)=\frac{1}{\prod_{k=0}^{K}a^{k}(2\uppi)^{K}|\Sigma|^{\frac{1}{2}}}e^{-\frac{1}{2}\left(h^{-1}(a)-\mu\right)^{T}\Sigma^{-1}\left(h^{-1}(a)-\mu\right)};\quad\text{for }a\in\operatorname{\mathbb{S}}^{K};$ (6)
see, e.g. Atchison and Shen (1980) for a detailed overview of the logistic-normal distribution.
We choose a policy that has logistic-normal distribution with density $\pi_{\theta}(\cdot\mid s)$ for each state $s.$ The parameters $\theta$ are the weights of a neural network (or any other parametric family of functions) that maps states $s$ to the mean of the underlying normal distribution $\mu.$ We could also make the algorithm learn the covariance matrix $\Sigma.$ But here, we only learn $\mu$ and model $\Sigma$ as $\sigma_{i}I_{K},$ where $I_{K}$ is the $K$-dimensional identity matrix and $\sigma_{i}$ a constant that is reduced after each training epoch $i.$ This allows for exploration at the beginning of training but forces the algorithm to gradually reduce the variance of the stochastic policy until it becomes almost deterministic. Let $\varphi_{\theta}$ be the density of the underlying normal distribution on $\operatorname{\mathbb{R}}^{K}$ with mean $\mu_{\theta}$ and covariance matrix $\Sigma.$ Since the normalizing constant in (6) does not depend on $\theta,$ we obtain for a state $s$ and action $a\in\operatorname{\mathbb{S}}^{K}$ that
@@
$\nabla_{\theta}J(\pi_{\theta})$ $=\underset{a\sim\pi_{\theta}}{\operatorname{\mathbb{E}}}\left[\sum_{n=0}^{N-1}\nabla_{\theta}\log\left(\varphi_{\theta}\left(h^{-1}(a_{n})|s_{n}\right)\right)A^{\pi_{\theta}}(s_{n},a_{n})\right]$ (8)
$=\underset{x\sim\varphi_{\theta}}{\operatorname{\mathbb{E}}}\left[\sum_{n=0}^{N-1}\nabla_{\theta}\log\left(\varphi_{\theta}(x_{n}\mid s_{n})\right)A^{\pi_{\theta}}(s_{n},h(x_{n}))\right].$ (9)
Here $x\sim\varphi_{\theta}$ means that each element $x_{n}$ in the sequence $x=(x_{0},x_{1},\dots,x_{N-1})$ is sampled from a normal distribution with conditional density $\varphi_{\theta}(\cdot\mid s_{n})$ for each state $s_{n}$ in the sequence $(s_{0},s_{1},\dots,s_{N-1}).$ We emphasize that the gradient in (9) only affects the density of the underlying normal distribution. The elements $x_{n}\in\operatorname{\mathbb{R}}^{K}$ are then transformed with $h$ to obtain actions $a_{n}\in\operatorname{\mathbb{S}}^{K}$ without affecting the gradient.
@@
Instead of using the multivariate logistic-normal distribution, one could combine the multivariate normal distribution with a softmax activation function. More precisely, if $X$ is a multivariate normal distribution on $\operatorname{\mathbb{R}}^{K+1}$ we could parametrize the policy as $\text{softmax}(X)$, where $\text{softmax}(\cdot)$ maps from $x\in\operatorname{\mathbb{R}}^{K+1}$ to $a\in\operatorname{\mathbb{S}}^{K}$ with
@@
However, $\text{softmax}(X)$ is not bijective. Therefore, it is not possible to derive the policy gradient in closed form as in (9).
@@ 4.3 Initialization and scaling of policy parameters
The initialization of the mean and the scaling of the variances of the stochastic policy $\pi_{\theta}$ is an important aspect of training the policy. The mean and covariance of the logistic-normal distribution are not known in closed form. However, for the log ratios, we have
@@
for $j,k,l,m\in\{0,1,2,\dots K\},$ with the convention $\mu^{K}=0$ and $\sigma^{j,K}=0.$ We set the parameters of the last layer of the neural network corresponding to the policy to zero at initialization with a bias term $b\in\operatorname{\mathbb{R}}^{K}.$ Then, the bias $b$ controls the initial allocations of the policy. Furthermore, we choose a diagonal covariance matrix to ensure that unrelated log ratios are uncorrelated. Indeed, for different $j,k,l$ and $m,$ equation (11) implies
$\operatorname{\mathrm{Cov}}\left(\log\left(\frac{a^{j}}{a^{k}}\right),\log\left(\frac{a^{l}}{a^{m}}\right)\right)=0,$
@@
$\operatorname{\mathrm{Var}}\left(\log\left(\frac{a^{j}}{a^{k}}\right)\right)=\begin{cases}\sigma^{jj}:\text{ if }j=K\text{ or }k=K,\\ \sigma^{jj}+\sigma^{kk}:\text{ otherwise}.\end{cases}$
The initial parameters of the covariance matrix $\Sigma$ control the initial exploration rate of the policy. As training progresses, we gradually lower the entries in the covariance matrix to discourage exploration and encourage exploitation. For $H\in\operatorname{\mathbb{N}}$ gradient steps and for an initial variance $\sigma_{\text{init}}>0$ and final variance $\sigma_{\text{final}}>0,$ we set the variances on the diagonal at each gradient step $i\in\{1,2,\dots,H\}$ to
$\sigma^{kk}_{i}=(\sigma_{\text{final}}-\sigma_{\text{init}})\left(\frac{i-1}{H-1}\right)+\sigma_{\text{init}},\quad k=0,1\dots,K-1$ (12)
The choice of the initial and final variances depends on the application.
@@ 4.4 Empirical policy gradient and advantage function estimation
In practice, the expectation (9) is not available, and we replace it with a sample-based estimate. This also requires an estimate of the advantage function (5). We estimate it using an estimate of the value function $V_{\vartheta}^{\pi_{\theta}},$ where $\vartheta$ are the weights of a neural network. We make $H\in\operatorname{\mathbb{N}}$ gradient steps with respect to the policy parameters $\theta$ and the parameters $\vartheta$ corresponding to the value function. At the beginning of the training cycle, the parameters are initialized as $\theta_{1}$ and $\vartheta_{1}.$ In each training epoch $i\in\{1,\dots,H\},$ we collect $\tau\in\operatorname{\mathbb{N}}$ trajectories
@@
where the actions $a_{n,k}$ are generated by sampling $x_{n,k}$ from the underlying normal distribution with density $\varphi_{\theta_{i}}(\cdot\mid s_{n,k})$ and transformed with $h.$ For each state action pair $(s_{n,k},a_{n,k}),$ the advantage function is estimated as
@@
The parameters $\theta$ of the policy neural network are updated, from $\theta_{i}$ to $\theta_{i+1},$ by making a gradient step with learning rate $\eta>0$ using the loss function
$\theta\mapsto-\frac{1}{\tau N}\sum_{k=1}^{\tau}\sum_{n=0}^{N-1}\log\varphi_{\theta}(x_{n,k}\mid s_{n,k})A_{\vartheta_{i}}^{\pi_{\theta_{i}}}(s_{n,k},a_{n,k}).$ (15)
@@
Algorithm 1 Actor-critic policy gradient with logistic-normal
@@
     Collect trajectories $\mathcal{T}=\left\{(s_{n,k},a_{n,k},r(s_{n,k},a_{n,k}),n\in\{0,\dots,N\},k\in\{1,\dots,\tau\}\right\},$ by sampling $x_{n,k}\sim\varphi_{\theta},$ and transforming $a_{n,k}=h(x_{n,k}).$
@@
Schulman et al. (2015) describe a general method for estimating advantage functions that interpolates between high-variance/low-bias estimates and low-variance/high-bias estimates. Estimate (14) has low bias but high variance. We reduce the variance of the estimate by using a large number of trajectories $\tau$ at each gradient step.
@@ 5 Market environment
We test our RL algorithm in a market environment populated by three different types of traders: noise traders who place and cancel orders randomly, tactical traders who react to order book imbalance, and a strategic trader who buys or liquidates a large position with a TWAP strategy, and therefore, trades in one direction at constant speed. We simulate the order book in the interval $[-\Delta t,T].$
@@ 5.1 Noise traders
Our first category of traders are noise traders who submit market, limit, and cancellation orders according to independent Poisson processes (Cont et al., 2010; Abergel and Jedidi, 2013). The intensities of market orders, limit orders, and cancellations arrivals are modeled in the following way.
@@
  • For $k\in\{1,\dots,D\},$ limit buy and sell orders arrive with intensity $\lambda^{L,k}$ at $k$ ticks below the best ask price and $k$ ticks above the best bid price, respectively.
  • If the spread is $j$ ticks, then for $k\in\{j,\dots,D\},$ cancellations of limit buy orders arrive with intensity $\lambda^{C,k}v^{b,{k-j+1}}$ at $k$ ticks below the best ask price, and cancellations of limit sell orders arrive with intensity $\lambda^{C,k}v^{a,{k-j+1}}$ at $k$ ticks above the best bid price.

### v2

@@ 3.4 Benchmark algorithms
We compare our trained reinforcement learning algorithm with two heuristic benchmarks and an alternative reinforcement learning algorithm using the Dirichlet distribution. The Dirichlet algorithm is explained in Appendix B.2. The heuristic strategies trade according to the following rules:
@@
  • Time-weighed average price [^3] This is abbreviated with TWAP and represents a standard execution algorithm banks and hedge funds use. algorithm (TWAP algorithm): Divides the position into blocks of size $M/N$ and posts a limit sell order of size $M/N$ at each time $t_{n},n=0,\dots,N-1.$
@@ 4.1 Actor-critic policy gradient
The trade execution problem described in Section 3 is a stochastic control problem. The transition probabilities from state $s_{n}$ to $s_{n+1}$ and the reward function $r(\cdot)$ are not available in closed form, but we can simulate them. Therefore, the problem can be solved with reinforcement learning. Actor-critic policy gradient algorithms are a class of reinforcement learning algorithms that aim to find an optimal stochastic policy $\pi_{\theta}(\cdot\mid s_{n}),$ referred to as actor, by estimating the gradient of the objective function $J(\pi_{\theta}).$ A separate parameterized value or advantage function, called critic, is used to reduce the variance of the gradient estimate. We use a neural network with weights $\theta$ to map states $s_{n}$ to the parameters of the policy’s density. The gradient of the objective function can be computed as
$\nabla_{\theta}J(\pi_{\theta})=\underset{\begin{subarray}{c}a\sim\pi_{\theta}\\ s_{0}\sim\rho\end{subarray}}{\operatorname{\mathbb{E}}}\left[\sum_{n=0}^{N-1}A^{\pi_{\theta}}(s_{n},a_{n})\nabla_{\theta}\log\pi_{\theta}(a_{n}\mid s_{n})\right],$ (4)
by the policy gradient theorem; see e.g. Schulman et al. (2015). We note that the actions $a_{0},\dots a_{N-1}$ are sampled from a conditional distribution with density $\pi_{\theta}(\cdot\mid s_{n})$ and the states $(s_{0},\dots,s_{N-1})$ are sampled from the market environment. Here, $A^{\pi_{\theta}}$ is the advantage function, which is defined together with the value function $V^{\pi_{\theta}}$ by
$V^{\pi_{\theta}}(s)$ $=\underset{a\sim\pi_{\theta}}{\operatorname{\mathbb{E}}}\left[\sum_{l=n}^{N-1}r(s_{l},a_{l})\mid s_{n}=s\right],$ (5)
$A^{\pi_{\theta}}(s,a)$ $=\underset{a\sim\pi_{\theta}}{\operatorname{\mathbb{E}}}\left[\sum_{l=n}^{N-1}r(s_{l},a_{l})\mid s_{n}=s,a_{n}=a\right]-V^{\pi_{\theta}}(s),$ (6)
@@ 4.2 Policy gradient with logistic-normal distribution
In most applications of policy gradients with continuous action spaces, the multivariate normal distribution is used as an action distribution, see e.g. Schulman et al. (2017). However, in our application, the actions must be contained in the simplex. Therefore, a natural choice for the action distribution is the logistic-normal distribution, which has support in $\operatorname{\mathbb{S}}^{K}.$
@@
$\pi(a)=\frac{1}{(2\uppi)^{K}|\Sigma|^{\frac{1}{2}}\prod_{k=0}^{K}a^{k}}e^{-\frac{1}{2}\left(h^{-1}(a)-\mu\right)^{T}\Sigma^{-1}\left(h^{-1}(a)-\mu\right)},\quad\text{for }a\in\operatorname{\mathbb{S}}^{K};$ (7)
see e.g. Atchison and Shen (1980) for a detailed overview of the logistic-normal distribution.
We choose a policy that has logistic-normal distribution with density $\pi_{\theta}(\cdot\mid s)$ for each state $s.$ The parameters $\theta$ are the weights of a neural network (or any other parametric family of functions) that maps states $s$ to the mean of the underlying normal distribution $\mu.$ We model the covariance matrix $\Sigma$ as a hyperparameter that is iteratively reduced during training, as explained further below. Let $\varphi_{\theta}$ be the density of the underlying normal distribution on $\operatorname{\mathbb{R}}^{K}$ with mean $\mu_{\theta}$ and covariance matrix $\Sigma.$ Since the normalizing constant in (7) does not depend on $\theta,$ we obtain for each state $s$ and action $a\in\operatorname{\mathbb{S}}^{K}$ that
@@
$\nabla_{\theta}J(\pi_{\theta})$ $=\underset{\begin{subarray}{c}a\sim\pi_{\theta}\\ s_{0}\sim\rho\end{subarray}}{\operatorname{\mathbb{E}}}\left[\sum_{n=0}^{N-1}A^{\pi_{\theta}}(s_{n},a_{n})\nabla_{\theta}\log\left(\varphi_{\theta}\left(h^{-1}(a_{n})|s_{n}\right)\right)\right]$ (9)
$=\underset{\begin{subarray}{c}x\sim\varphi_{\theta}\\ s_{0}\sim\rho\end{subarray}}{\operatorname{\mathbb{E}}}\left[\sum_{n=0}^{N-1}A^{\pi_{\theta}}(s_{n},h(x_{n}))\nabla_{\theta}\log\left(\varphi_{\theta}(x_{n}\mid s_{n})\right)\right].$ (10)
Here $x\sim\varphi_{\theta}$ means that each element $x_{n}$ in the sequence $x=(x_{0},x_{1},\dots,x_{N-1})$ is sampled from a normal distribution with conditional density $\varphi_{\theta}(\cdot\mid s_{n})$ for each state $s_{n}$ in the sequence $(s_{0},s_{1},\dots,s_{N-1}).$ The main difference between our algorithm and other actor-critic reinforcement learning algorithms is the use of the logistic-normal distribution, which ensures actions lie in the simplex. The representation (10) still allows us to compute gradients efficiently.
@@
Instead of using the multivariate logistic-normal distribution, one could combine the multivariate normal distribution with a softmax activation function. More precisely, if $X$ is a multivariate normal distribution on $\operatorname{\mathbb{R}}^{K+1}$ we could parametrize the policy as $\text{softmax}(X)$, where $\text{softmax}(\cdot)$ maps $x\in\operatorname{\mathbb{R}}^{K+1}$ to $a\in\operatorname{\mathbb{S}}^{K}$ with
@@
However, $\text{softmax}(X)$ is not bijective. Therefore, it is not possible to derive the policy gradient in closed form as in (10). Prior works such as Tian et al. (2022) and Winkel et al. (2023) have employed the Dirichlet distribution to model actions on the simplex. We use an actor-critic algorithm based on the Dirichlet distribution as a benchmark in our experiments, but we find that it performs sub-optimally, even with extensive hyperparameter tuning.
@@ 4.3 Initialization and scaling of policy parameters
The initialization of the mean and reduction of the variances of the stochastic policy $\pi_{\theta}$ is an important aspect of training the policy. The mean and covariance of the logistic-normal distribution are not known in closed form. However, for the log ratios, we have
@@
for $j,k,l,m\in\{0,1,2,\dots K\},$ with the convention $\mu^{K}=0$ and $\Sigma^{j,K}=0.$ At initialization, we set the parameters of the neural network’s final policy layer close to zero, except for the bias term $b\in\operatorname{\mathbb{R}}^{K}.$ Then, the bias $b$ controls the initial allocations of the policy. We model $\Sigma$ as $\sigma I_{K},$ where $I_{K}$ is the $K$-dimensional identity matrix and $\sigma$ is reduced after each gradient step. This allows for exploration at the beginning of training but forces the algorithm to gradually reduce the variance of the stochastic policy until it becomes deterministic. With a diagonal covariance matrix, for different $j,k,l$ and $m,$ equation (12) implies
$\operatorname{\mathrm{Cov}}\left(\log\left(\frac{a^{j}}{a^{k}}\right),\log\left(\frac{a^{l}}{a^{m}}\right)\right)=0.$
@@
$\operatorname{\mathrm{Var}}\left(\log\left(\frac{a^{j}}{a^{k}}\right)\right)=\begin{cases}\sigma:\text{ if }j=K\text{ or }k=K,\\ 2\sigma:\text{ otherwise}.\end{cases}$
The initial parameters of the covariance matrix $\sigma I_{K}$ determine the policy’s initial exploration rate. As training progresses, we gradually lower the entries in the covariance matrix to discourage exploration and encourage exploitation. For $H\in\operatorname{\mathbb{N}}$ gradient steps and for an initial variance $\sigma_{\text{init}}>0$ and final variance $\sigma_{\text{final}}>0,$ we set the covariance matrix to $\sigma_{i}I_{K},$ where
$\sigma_{i}=(\sigma_{\text{final}}-\sigma_{\text{init}})\left(\frac{i-1}{H-1}\right)+\sigma_{\text{init}},\quad\text{for }i=1,2,\dots,H.$ (13)
The choice for the initial and final variance and the bias term for the trade execution problem is discussed further below. For other applications, they should be chosen so that exploration is encouraged at the beginning and discouraged at the end of the training cycle. For example, in optimal portfolio allocation problems, one should set the bias such that the initial allocation is balanced to all assets in the portfolio, ensuring that the algorithm observes a complete set of market scenarios.
Remark 4.2.
Rather than manually scaling the algorithm’s action variances as in (13), they can also be encoded and learned during training. We explore this option in Appendix C.2, where we model the covariance matrix as a learnable diagonal matrix. We find that manual scaling performs significantly better in one of our test cases, but similarly in all other test cases. Although learning a full (non-diagonal) covariance matrix is also possible, we expect this to complicate the learning process and do not pursue it here.
@@ 4.4 Empirical policy gradient and advantage function estimation
In practice, the expectation (10) is not available, but can be replaced with a sample-based estimate. This also requires an estimate of the advantage function (6). We estimate it using an estimate of the value function $V_{\vartheta}^{\pi_{\theta}},$ defined in (5), where $\vartheta$ are the weights of a neural network. We make $H\in\operatorname{\mathbb{N}}$ gradient steps with respect to the policy parameters $\theta$ and the parameters $\vartheta$ corresponding to the value function. At the beginning of the training cycle, the parameters are initialized as $\theta_{1}$ and $\vartheta_{1}.$ In each training epoch $i\in\{1,\dots,H\},$ we collect $\tau\in\operatorname{\mathbb{N}}$ trajectories
@@
where the actions $a_{n,k}$ are generated by sampling $x_{n,k}$ from the underlying normal distribution with density $\varphi_{\theta_{i}}(\cdot\mid s_{n,k}),$ which are then transformed with $h.$ For each state-action pair $(s_{n,k},a_{n,k}),$ the advantage function is estimated as
@@
The parameters $\theta$ of the policy network are updated, from $\theta_{i}$ to $\theta_{i+1},$ by making a gradient step with learning rate $\eta>0$ using the loss function
$\theta\mapsto-\frac{1}{\tau N}\sum_{k=1}^{\tau}\sum_{n=0}^{N-1}A_{\vartheta_{i}}^{\pi_{\theta_{i}}}(s_{n,k},a_{n,k})\log\varphi_{\theta}(x_{n,k}\mid s_{n,k}).$ (16)
@@
Algorithm 1 Actor-critic policy gradient method with logistic-normal
@@
  Collect trajectories $\mathcal{T}=\left\{(s_{n,k},a_{n,k},r(s_{n,k},a_{n,k})),n\in\{0,\dots,N\},k\in\{1,\dots,\tau\}\right\},$ by sampling $x_{n,k}\sim\varphi_{\theta},$ and transforming $a_{n,k}=h(x_{n,k}).$
@@
Schulman et al. (2015) describe a general method, which they call generalized advantage function estimation (GAE), for estimating advantage functions that interpolates between high-variance/low-bias estimates and low-variance/high-bias estimates. The interpolation is controlled with a parameter [^4] We use the symbol $\lambda^{\text{GAE}}$ rather than $\lambda,$ as in their paper, to avoid confusion with the symbol for trading intensities below. $\lambda^{\text{GAE}}\in[0,1].$ In (15), we choose a high-variance/low-bias estimate, by setting $\lambda^{\text{GAE}}=1$ and reduce the variance of the estimate by using a large number of trajectories $\tau$ at each gradient step. The hyperparameters of the algorithm are discussed in B.1. For comparison, we conduct an additional experiment with a smaller number of trajectories per gradient step and different choices of $\lambda^{\text{GAE}},$ to reduce variance, in Appendix C.1. We find that choosing a smaller value for $\lambda^{\text{GAE}}$ does not lead to better results.
@@ 5 Market environment
We test our RL algorithm in a market environment populated by three different types of traders: noise traders who place and cancel orders randomly, tactical traders who react to order book imbalance, and a strategic trader who buys or liquidates a large position with a TWAP strategy, and therefore, trades in one direction at constant speed. We simulate the order book during the time interval $\Delta t,T].$
@@ 5.1 Noise traders
Our first category of traders are noise traders who submit market, limit, and cancellation orders according to independent Poisson processes (Cont et al., 2010; Abergel and Jedidi, 2013). The intensities of market orders, limit orders, and cancellations are modeled in the following way.
@@
  • For $k\in\{1,\dots,D\},$ limit buy and sell orders arrive $k$ ticks below the best ask price or $k$ ticks above the best bid price with intensity $\lambda^{L,k}.$
  • If the spread is $j$ ticks, then for $k\in\{j,\dots,D\},$ cancellations of limit buy orders arrive $k$ ticks below the best ask price with intensity $\lambda^{C,k}v^{b,{k-j+1}}$, and cancellations of limit sell orders arrive $k$ ticks above the best bid price with intensity with intensity $\lambda^{C,k}v^{a,{k-j+1}}.$